OpenStudy (anonymous):

(A/b) + (b/a) / (1/a) - (1/b). Help needed

OpenStudy (tkhunny):

Did you mean this? \(\dfrac{\dfrac{a}{b}+\dfrac{b}{a}}{\dfrac{1}{a}-\dfrac{1}{b}}\) If so, that is not very close to what you have written.